use multiplication principle to solve equation;n x 4/15= 12/15
Answer by stanbon(57377)

use multiplication principle to solve equation;
n x 4/15= 12/15
---
Multiply both sides by 15/4 to get:
n = (12/15)(15/4)
---
n = 3
======================
Cheers,
Stan H.